OpenStudy (anonymous):

Could some one check my answers? (Algebra 1, Functions) Will give medal.

OpenStudy (anonymous):

what school do you go to

OpenStudy (anonymous):

FLVS why?

OpenStudy (ccswims):

Omg me too

OpenStudy (ccswims):

I don't feel alone now

OpenStudy (anonymous):

haha.

OpenStudy (anonymous):

just seeing if i did the quiz already

OpenStudy (anonymous):

so what are you questions

OpenStudy (anonymous):

The table below shows the values of y for each value of x: Part A: Does the table represent a relation that is a function? Justify your answer by using the values shown in the table. (4 points) Part B: The function f(x) shown below represents the number of baskets Jess made at different distances, x, in meters, from the hoop: f(x) = -x + 6 Calculate and interpret the meaning of f(4). (4 points) Part C: Write an ordered pair to represent the input and output of the function in Part B when Jess is at a distance of p meters from the hoop. (2 points)

OpenStudy (amistre64):

passing answers back and forth will not be tolerated ... just saying.

OpenStudy (anonymous):

For the part a: I said This table is not a function, because the x values have more than one output. ex: (1,0) (1,-1) Part B: : f(4)=-4+6=2 this means that Jess made 4 baskets from 2 meters from the hoop. Part C: (-4,2)

OpenStudy (anonymous):

no i would not just give him the answers i would just be able to help him more

OpenStudy (anonymous):

I don't want answers I just want to know how I can correct my self if something is wrong. @amistre64

OpenStudy (amistre64):

just making sure :)

OpenStudy (jdoe0001):

A) is correct, there is one x-repeat, thus not a function B) f(x) or "y" represents the baskets made by Jess f(4) means she made only 2 baskets, y=-4+6 = 2 from 4 meters, or "x" C) see above

OpenStudy (anonymous):

ok I see what I did wrong for part B, so should the numbers be switched around for part c?

OpenStudy (jdoe0001):

well... actually... .just noticed... part C) uses "p" meters so \(\bf y=-4{\color{brown}{ x}}+6\qquad x={\color{brown}{ p}}\qquad y=?\)

OpenStudy (anonymous):

I have no idea how to find y without finding p.

OpenStudy (jdoe0001):

well... don't find "p" then... leave it in "p" terms

OpenStudy (anonymous):

(p,-4p+6) ?

OpenStudy (jdoe0001):

(p,-4p+6) \(\Large \checkmark\)

OpenStudy (anonymous):

but shouldn't I answer the question with actual numbers?

OpenStudy (jdoe0001):

well. yes.. IF the input were a number but in this case is just a variable

OpenStudy (anonymous):

ok thanks so much!

OpenStudy (jdoe0001):

yw